OpenStudy (anonymous):

Fe-61 has a half life of 6.00 min. Of the 100.0 mg sample, how much will remain after 18.0 min?

OpenStudy (anonymous):

\[\frac{1}{2}=e^{6.00k}~~\implies~~k=\frac{1}{6}\ln\frac{1}{2}\approx -0.1155\] After 18 minutes: \[y=100e^{18\times k}=12.5\]